OpenStudy (anonymous):

2sinx-2sinxcosx^2x+1-cos^2x

myininaya (myininaya):

what are we doing?

OpenStudy (akshay_budhkar):

i think he wants it simplified o.O

myininaya (myininaya):

I can rewrite it I don't know if it is considered simplified \[2\sin(x)-2\sin(x) \cos^2(x)+1-\cos^2(x)=2\sin(x)(1-\cos^2(x))+1-\cos^2(x)\] \[=2 \sin(x) (1-\cos^2(x))+1(1-\cos^2(x))=(1-\cos^2(x))(2 \sin(x)+1)\] \[\sin^2(x)(2 \sin(x)+1)\]

OpenStudy (akshay_budhkar):

umm yea looks good myininaya.. or maybe that stuff is 0..so we can find x i believe

myininaya (myininaya):

good esp

myininaya (myininaya):

if so set both factors =0

OpenStudy (akshay_budhkar):

yea :D

OpenStudy (anonymous):

It says "Solve each equation for \[0 \le x < 360\]

OpenStudy (akshay_budhkar):

it is not an equation <.<

myininaya (myininaya):

just so you know you didn't write in equation

myininaya (myininaya):

an*

OpenStudy (akshay_budhkar):

it is an expression

myininaya (myininaya):

so it is =0 like akshay said?

OpenStudy (anonymous):

yes, it is equal to 0

myininaya (myininaya):

so i factored your expression set both of the factors in my expression =0

OpenStudy (akshay_budhkar):

can you do that?

myininaya (myininaya):

a*b=0 => a=0 or b=0 or both=0

OpenStudy (akshay_budhkar):

Habu can u proceed further? or do u want it to be done?

OpenStudy (anonymous):

Can you write it all out please and thank you (:

myininaya (myininaya):

\[\sin^2(x)(2 \sin(x)+1)=0 => \sin^2(x)=0 \text{ or } 2 \sin(x)+1=0 \]

myininaya (myininaya):

you have two equations to solve

myininaya (myininaya):

sin(x)=0 is the easiest

myininaya (myininaya):

sin(0)=0 right so x=0 is one solution now you try to solve 2sin(x)+1=0

myininaya (myininaya):

i will give you a clue subtract one on both sides then divide 2 on both sides

OpenStudy (akshay_budhkar):

x=180 is another soluton of sinx=0 :)

myininaya (myininaya):

oh yes and that one

OpenStudy (akshay_budhkar):

will do the second one? @myininaya you r too quick lol :P

OpenStudy (anonymous):

7pie/6 and 11pie/6

OpenStudy (akshay_budhkar):

yes correct :)

OpenStudy (anonymous):

Thank you both for your help (:

OpenStudy (akshay_budhkar):

yw :D.. lol thank myininaya i did nothing :D

myininaya (myininaya):

 i would write my answer in degrees since your inequality is written in degrees